Please help, I don't understand how to find out their individual weight!
4 lb of peanuts and 3 lb of cashews cost $16.50, while 5 lb of peanuts and 7 lb of cashews cost $32.00. Find the cost per lb of each.

Answer by scott8148(6628)

the system of equations is

4p + 3c = 16.50

5p + 7c = 32.00

solving by elimination

1st eqn times -5 ___ -20p - 15c = -82.50

2nd eqn times 4 ___ 20p + 28c = 128.00

adding eqns (to eliminate p) ___ 13c = 45.50 ___ c = 3.50

substituting (to find p) ___ 4p + 3(3.50) = 16.50 ___ 4p = 6.00 ___ p = 1.50